### v1

# A Minimal Subset Approach for Efficient and Scalable Loop Closure

Abstract
Loop closure detection in large-scale and long-term missions can be computationally demanding due to the need to identify, verify, and process numerous candidate pairs to establish edge connections for the pose graph optimization. Keyframe sampling mitigates this by reducing the number of frames stored and processed in the back-end system. In this article, we address the gap in optimized keyframe sampling for the combined problem of pose graph optimization and loop closure detection. Our Minimal Subset Approach (MSA) employs an optimization strategy with two key factors, redundancy minimization and information preservation, within a sliding window framework to efficiently reduce redundant keyframes, while preserving essential information. This method delivers comparable performance to baseline approaches, while enhancing scalability and reducing computational overhead. Finally, we evaluate MSA on relevant publicly available datasets, showcasing that it consistently performs across a wide range of environments, without requiring any manual parameter tuning.
Index Terms:
Loop Closure, Place Recognition, Localization, Keyframe Sampling, Optimization.
I Introduction
Simultaneous Localization and Mapping (SLAM) in large-scale and long-term missions presents significant challenges, particularly in maintaining accuracy and efficiency as the environment grows over time. A key component of SLAM is Pose Graph Optimization (PGO), which ensures consistent mapping by optimizing the global trajectory. However, this process relies heavily on loop closure detection, to reduce accumulated drift, which is computationally expensive, especially in large-scale environments, due to the vast number of candidate pairs that need to be identified, verified, and processed to establish node connections in the pose graph. As the mission progresses, the number of loop closure candidates grows rapidly, leading to significant delays in processing times, with some datasets generating over 150,000 candidates, requiring four and a half hours to process, exceeding that of the mission time [1]. This complexity is further compounded by the computational intensity of data association and sequence matching during map optimization, where the processing time can range from seconds to several minutes, depending on the size of the data [2, 3].
Due to these computational challenges, many SLAM systems in large-scale environments have opted to either avoid or minimize loop closure detection to reduce the burden on limited computational resources. Some approaches, particularly in expansive environments, forego loop closures entirely, as the growing number of candidates becomes increasingly difficult to handle in real-time applications, as highlighted recently in [4]. While loop closure detection is critical for correcting drift and ensuring map consistency, the computational complexity can overwhelm systems if not managed effectively, often leading to false positives that degrade the overall quality of the map [5].
I-A Related Work
Efficient implementations of pose graph optimization, such as GTSAM [6], g2o [7] and iSAM2 [8] have significantly enhanced its scalability by leveraging factor graphs and incremental optimization techniques. Additionally, sparsification methods [9, 10, 11] and hierarchical approaches [12, 13] have further improved computational efficiency by reducing the number of active nodes in the optimization process. However, despite these advances, loop closure detection remains a major bottleneck due to the large number of candidate pairs that must be checked and processed to add meaningful constraints to the pose graph.
Figure 1: An example on KITTI 06. Comparison of two sampled trajectories: one using an entropy-based approach and the other using the proposed Minimal Subset Approach (MSA). The MSA-sampled trajectory achieves a lower Relative Pose Error (RPE) after the pose graph optimization.
Keyframe sampling helps reduce the computational demands of loop closure detection by limiting the number of frames stored and processed in SLAM systems. In LiDAR-(Inertial) Odometry, systems such as LIO-SAM [14, 15, 16] or back-end global optimization systems like LAMP [17, 18], use fixed Euclidean distance intervals (1–2 meters) for keyframe generation to maintain local maps and optimize global poses. Although effective, these methods lack adaptability in dynamic environments with varying densities of loop closure candidates. Recent advancements have introduced adaptive keyframe sampling techniques, adjusting intervals based on environmental spaciousness [19, 20], although manual threshold adjustment is still required. Entropy-based methods, like the one proposed by [21], use information theory to select keyframes, but face similar adaptability issues. Other approaches, such as [22], focus on displacement vector similarity to balance computational cost and map completeness while managing keyframes in a sliding window.
In this article, we address the gap in optimized keyframe sampling for pose graph optimization and loop closure detection by streamlining keyframe extraction. Our Minimal Subset Approach uses two key criteria, redundancy minimization and information preservation, within a sliding window framework to efficiently reduce redundant keyframes while maintaining essential information. This method enhances adaptability across various environments and eliminates the need for manual threshold tuning, with an example of its performance illustrated on Fig. 1.
I-B Contributions
Based on the aforementioned, the contributions of this article can be summarized as follows: (a) We propose a Minimal Subset Approach (MSA) that efficiently addresses the problem of loop closure detection and pose graph optimization by leveraging a subset of keyframes. This method achieves comparable results to baseline approaches but with improved scalability and efficiency, reducing computational demands. (b) The proposed approach is environment agnostic, avoiding manual parameter tuning and offering consistent performance across diverse environments. This is achieved through a sliding window optimization in the hyper-dimensional descriptor space, eliminating the need for threshold adjustments. (c) The MSA addresses the combined problem of pose graph optimization and loop closure detection. Unlike many existing methods that focus solely on extracting keyframes for scan matching and pose graph optimization, MSA takes into account how sampling intervals impact loop closure detection performance through place recognition, providing a more integrated solution to these interrelated challenges.
II Problem Formulation
II-A Pose Graph Optimization
The objective of PGO is to find the most likely configuration of a robot’s trajectory by minimizing the error in a graph-based representation of the poses. The pose graph is represented as a set of nodes and edges, where each node represents a robot pose and each edge encodes a spatial constraint between two poses [23]. Let $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ be the set of $N$ robot poses, where, $\mathbf{x}\in SE(3)$, $\mathbf{z}_{ij}=h(\mathbf{x}_{i},\mathbf{x}_{j})$ be the observed spatial constraint between a pair of poses $\langle i,j\rangle$ and $\hat{\mathbf{z}}_{ij}(\mathbf{x}_{i},\mathbf{x}_{j})$ be the predicted value of that constraint. The error $\mathbf{e}_{ij}\in\mathbb{R}^{6}$ quantifies the difference between the observed relative transformation and the predicted transformation between poses $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ and is denoted as:
$\mathbf{e}_{ij}=\mathbf{e}(\mathbf{x}_{i},\mathbf{x}_{j})=\mathbf{z}_{ij}-\hat{\mathbf{z}}_{ij}(\mathbf{x}_{i},\mathbf{x}_{j})$ (1)
We can then define the set of edges =$\mathcal{E}=\big{\{}\langle i,j\rangle\>|\>\exists\,\mathbf{z}_{ij}\big{\}}$, that contains all the pairs in which a relative constraint exists. The goal of pose graph optimization is to find the optimal configuration of poses $\mathbf{X}^{*}$ that minimizes the overall error function $E(\mathbf{X})$, denoted as:
$\mathbf{X}^{*}=\operatorname*{arg\,min}_{\mathbf{X}}\sum_{\langle i,j\rangle\in\mathcal{E}}\mathbf{e}(\mathbf{x}_{i},\mathbf{x}_{j})^{\top}\,\mathbf{\Omega}_{ij}\,\mathbf{e}(\mathbf{x}_{i},\mathbf{x}_{j}),$ (2)
where $\mathbf{\Omega}_{ij}\in\mathbb{R}^{6\times 6}$ is the information matrix that reflects the certainty of the relative pose measurement for each pair.
II-B Odometry and Loop Closure Edges
Pose graph optimization is widely used as a global optimization method to mitigate accumulated pose drift during large-scale and long-term missions. This would not be possible without loop closures. Relying solely on pose-to-pose odometry constraints is insufficient to maintain global consistency, hence, loop closures are essential. Given this, it is reasonable to assume that the set of edges $\mathcal{E}$ in the pose graph can be partitioned into two distinct subsets. The first subset, $\mathcal{E}_{o}\subseteq\mathcal{E}$, consists of odometry edges, which represent consecutive movements of the robot based on local motion estimates, while the second subset, $\mathcal{E}_{l}\subseteq\mathcal{E}$, consists of loop closure edges, which correspond to observations of previously visited locations. Thus, the overall error function can be decomposed into two components:
$E(\mathcal{\mathbf{X}})=\sum_{\langle i,j\rangle\in\mathcal{E}_{o}}\mathbf{e}_{ij}^{\top}\mathbf{\Omega}_{ij}\,\mathbf{e}_{ij}+\sum_{\langle i,j\rangle\in\mathcal{E}_{l}}\mathbf{e}_{ij}^{\top}\mathbf{\Omega}_{ij}\,\mathbf{e}_{ij},$ (3)
where the edge sets satisfy the following two relationships: $\mathcal{E}=\mathcal{E}{o}\cup\mathcal{E}_{l}$ and $\mathcal{E}_{o}\cap\mathcal{E}_{l}=\varnothing$.
II-C Loop Closure Detection
Detecting loop closures is challenging, particularly in large-scale missions, where accumulated drift can cause traditional radius-based searches to fail by missing nearby candidates. A common solution is place recognition, which compares current sensory data (e.g., images or LiDAR scans) against a database of past observations to identify matches. To formalize this, we define a keyframe set $\mathbf{K}=\{(\mathbf{x}_{i},\mathbf{d}_{i},\mathbf{z}_{i})\,|\,i=1,\ldots,N\}$, where $\mathbf{x}_{i}\in SE(3)$ is the pose of keyframe $i$, $\mathbf{d}_{i}\in\mathbb{R}^{M}$ represents its descriptor which encodes distinct features [24, 25], and $\mathbf{z}_{i}$ is the set of associated measurements. The goal is to find loop closure candidates $\mathcal{L}$ by comparing descriptors:
=$\mathcal{L}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}\times\mathbf{K},\;\mathbf{k}_{i}\neq\mathbf{k}_{j}\big{\}},$ (4)
where $f(\mathbf{d}_{i},\mathbf{d}_{j})$ is a similarity function (e.g., cosine similarity, Euclidean distance) and $\tau$ is a threshold for considering two keyframes representing the same place. This approach generalizes to various scenarios, such as comparing a current keyframe with the $k$ nearest past keyframes in single-robot detection or comparing keyframes across multiple robots.
II-D Reducing Search Space in Place Recognition
The detection and verification of loop closure edges $\mathcal{E}_{l}\triangleq\mathcal{L}$ is computationally demanding due to the large number of potential keyframe pairs, which scales quadratically as $O(n^{2})$, where $n=|\mathbf{K}|$ is the number of keyframes. While using descriptors and methods like k-nearest neighbor (k-NN) search [26] can accelerate the process, additional steps are still required to remove outliers, mitigate false positives, or prioritize certain nodes [1], while extracting the relative pose transformation between loop closure nodes also involves computationally intensive techniques such as General Iterative Closest Point (GICP) [27] or Sample Consensus Initial Alignment (SAC-IA) [28]. This pipeline does not scale well for extensive missions, often resulting in processing times of 3–4 hours or more [4, 17, 18], and the challenge becomes even greater in multi-robot systems [2, 3].
To address this computational complexity, we propose a sampling strategy to reduce the number of keyframes considered for matching, while maintaining minimal impact on the loop closure set $\mathcal{L}$. Let $\mathbf{K}_{\text{S}}\subset\mathbf{K}$ represent a subset of keyframes selected through a sampling process $S:\mathbf{K}\rightarrow\mathbf{K}_{\text{S}}$ such that $|\mathbf{K}_{\text{S}}|\ll|\mathbf{K}|$. Our objective is to reduce the search space by confining the loop closure detection process to only the selected keyframes in $\mathbf{K}_{\text{S}}$. The goal is to identify a reduced set $\mathbf{K}_{\text{S}}^{*}$ that minimizes redundant keyframes while preserving the set of loop closures $\mathcal{L}$, a process denoted as:
$\mathbf{K}_{\text{S}}^{*}=\operatorname*{arg\,min}_{\mathbf{K}_{\text{S}}\subseteq\mathbf{K}}\left(|\mathbf{K}_{\text{S}}|\right),\;\text{subject to}\;\mathcal{L}^{*}=\mathcal{L},\text{ where }$ (5)
=$\mathcal{L}^{*}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}^{*}_{\text{S}}\times\mathbf{K}^{*}_{\text{S}}\big{\}},$ (6)
In this approach, the overall pose graph optimization is reformulated as follows:
$\mathbf{X}^{*}_{\text{S}}=\operatorname*{arg\,min}_{\mathbf{X}_{\text{S}}}\sum_{\langle i,j\rangle\in\mathcal{E}^{*}_{o}}\mathbf{e}_{ij}^{\top}\,\mathbf{\Omega}_{ij}\,\mathbf{e}_{ij}+\sum_{\langle i,j\rangle\in\mathcal{E}_{l}^{*}}\mathbf{e}_{ij}^{\top}\,\mathbf{\Omega}_{ij}\,\mathbf{e}_{ij},$ (7)
where $\mathbf{X}_{\text{S}}$ represents the poses corresponding to the optimally sampled keyframe set $\mathbf{K}_{\text{S}}^{*}$, and $\mathcal{E}_{o}^{*},\,\mathcal{E}_{l}^{*}\triangleq\mathcal{L}^{*}$ denote the odometry and loop closure edge pairs between the sampled keyframe poses $\mathbf{X}_{\text{S}}$, respectively.
II-E Underlying Challenges
Addressing the combined problem of pose graph optimization and loop closure detection through place recognition presents several underlying challenges that affect both the efficiency and accuracy of the solution.
II-E1 Impact of Sampling
While the sampling strategy $S$ aims to preserve key loop closure candidates, in practice it must retain a subset $\bar{\mathcal{L}}\subseteq\mathcal{L}$ that still ensures an effective solution for pose graph optimization. Redundant loop closure edges can lead to issues, such as overfitting, increased computational burden, and numerical instability in the optimization process. Moreover, because the sampling process reduces the set of poses, the odometry edge set $\mathcal{E}_{o}^{*}\subset\mathcal{E}_{o}$ is also reduced, which can affect the overall accuracy of the optimization  [29], with an example illustrated on Fig. 2.
II-E2 Combined Problem
Combining loop closure detection through place recognition with pose graph optimization presents challenges due to their differing characteristics. Loop closure detection relies on matching high-dimensional feature descriptors, which is computationally expensive and sensitive to environmental factors. Conversely, pose graph optimization relies on geometric consistency and is affected by noise and outliers in loop closures. Balancing these require careful consideration of both feature reliability and constraint robustness to prevent false positives and reduced efficiency from degrading optimization performance.
II-E3 Dynamic Keyframe Selection
To minimize the keyframe set size in real-time, we must consider the entire set $\mathbf{K}$; however, this is impractical for maintaining computational feasibility in loop closure detection. The challenge is to determine the contribution of each keyframe, which requires anticipating future keyframes and introduces a lack of casualty. Moreover, finding the optimal subset $\mathbf{K}_{\text{S}}^{*}$ is an NP-hard problem, involving an exhaustive search of all keyframe combinations, which becomes computationally infeasible for sets larger than 15–20 keyframes.
The goal of this research is to show that reducing the number of keyframes through an informed sampling strategy can significantly decrease the search space for loop closure detection, improving computational efficiency by reducing memory usage and query time without compromising SLAM accuracy. The proposed approach seeks to balance computational complexity and optimization precision, ensuring scalability and robustness.
Figure 2: Redundant Keyframes Effect. Example of removing redundant keyframes (light blue). The updated odometry edge set, $\mathcal{E}_{o}^{*}$, influences the pose graph optimization, while the sparsified loop closure edges, ¯$\bar{\mathcal{L}}=\mathcal{E}_{l}\backslash\langle 5,7\rangle$, can further reduce the computational complexity.
III Minimal Subset Approach
To address these challenges, we propose a Minimal Subset Approach (MSA) to approximate a solution to the problems in Eq. (5)-(6). The proposed MSA identifies and eliminates redundancy within a keyframe set while preserving crucial information. Designed for real-time implementation, it employs a sliding window combinatorial optimization with two distinct objectives, balancing information preservation and redundancy minimization.
III-A Redundancy in Keyframes
The frequency of a sensor, the platform’s speed, and the environmental characteristics can cause keyframe samples to capture redundant information if they are too closely spaced, with more details available on [30]. To address this, we define redundancy within a keyframe set and propose a metric to quantify it based on the descriptor space. A keyframe $\mathbf{k}$ is redundant within a set $\mathbf{K}$ if its removal does not affect the loop closure edge set $\mathcal{L}$ and does not create discontinuities in the map representation:
¯=$\bar{\mathcal{L}}\equiv\mathcal{L}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\bar{\mathbf{K}}\times\bar{\mathbf{K}}\big{\}},$
$\operatorname*{subject\>to}\>\>\>\delta_{l}\leqslant\|\mathbf{x}_{i}-\mathbf{x}_{i+1}\|_{2}\leqslant\delta_{u},\,\forall\,\mathbf{x}\in\mathbf{X},$ (8)
where $\bar{\mathbf{K}}=\mathbf{K}\,\backslash\{\mathbf{k}\}$. The spatial constraint in Eq. (8) ensures comprehensive map coverage and that consecutive odometry edges can be computed safely. The lower and upper pose distances, $\delta_{l}$ and $\delta_{u}$, usually range from 1 to 5 meters [24]. We quantify redundancy in a keyframe set using a metric that captures the similarity between consecutive keyframes, utilizing any similarity function $f_{\sigma}$, provided by the corresponding descriptor extraction framework:
$\rho_{\tau}(\mathbf{K})=\frac{1}{N-1}\sum_{i=1}^{N-1}f_{\sigma}(\mathbf{k}_{i},\mathbf{k}_{i+1}),$ (9)
where $0<\rho_{\tau}(\mathbf{K})\leqslant 1$ and $N$ is the number of keyframes in $\mathbf{K}$. Higher values indicate greater redundancy.
III-B Information Preservation in Keyframes
Let $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{M}$, represent the descriptor extraction function, either learning-based [24, 31] or handcrafted [25, 32], which maps each observation from 3D space to an $M$-dimensional representation. This function depends on the input observation (e.g. LiDAR scan), but it is understood that $\mathbf{z}_{t}\propto f(\mathbf{x}_{t})$ implies that each observation depends on the viewpoint and pose. To understand how the descriptors are sensitive to pose changes, we compute the Jacobian $\mathbf{J}$ of $F$ with respect to poses $\mathbf{x}$. Explicitly deriving the Jacobian can be impractical because of the intricate nature of functions like deep neural networks [30]. Consequently, we employ numerical approximations to estimate the rate of change, as shown in Fig. 3. Moreover, since poses are encompassed within $SE(3)$, we reduce the dimensional complexity by using the Euclidean norm for distance measurement instead of individual axis derivatives. It is essential to use yaw-invariant descriptors, as suggested in [24, 25, 33], to ensure the Jacobian calculations are meaningful by preventing orientation changes. Considering the descriptors $\mathbf{d}\in\mathbb{R}^{M}$ as random variables and poses $\mathbf{x}\in SE(3)$ as samples, the product $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ estimates the covariance matrix:
$\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}=\bigg{(}\frac{\partial\mathbf{F}}{\partial\mathbf{x}}\bigg{)}^{\mathsf{T}}\bigg{(}\frac{\partial\mathbf{F}}{\partial\mathbf{x}}\bigg{)}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$},$ (10)
where $\mathbf{\Lambda}$ is the diagonal matrix of eigenvalues and $\mathbf{V}$ is the matrix of eigenvectors from the decomposition. The eigenvectors represent the principal directions of variation, and the eigenvalues represent their magnitudes. The descriptors are transformed using the eigenvectors and scaled by the square root of the eigenvalues, $\mathbf{D}^{\prime}=\sqrt{\mathbf{\Lambda}}\cdot\mathbf{V}\cdot\mathbf{D}$, aligning them with the main directions of variability. Therefore, we can define the information preservation term for a keyframe set as:
$\pi_{\tau}(\mathbf{K})=-\frac{1}{N-1}\sum_{i=1}^{N-1}f_{\delta}(\mathbf{d}_{i}^{\prime},\mathbf{d}_{i+1}^{\prime}),$ (11)
where $-1\leqslant\pi_{\tau}(\mathbf{K})<0$, $f_{\delta}$ is the distance function between descriptors, and $\mathbf{d}^{\prime}\in\mathbf{D}^{\prime}$ are the transformed descriptors. Higher values of $\pi_{\tau}$ indicate that the poses better preserve the variability in the descriptor space. Eigenvectors $\mathbf{v}_{k}$ denote directions of maximal variability, and the $k$-th eigenvalue $\lambda_{k}$ quantifies the variance described by each eigenvector. Larger eigenvalues indicate more significant patterns of variability.
In summary, while both $\rho_{\tau}$ and $\pi_{\tau}$ utilize a similarity or distance function, they differ in objectives: the redundancy term focuses on local redundancy within the keyframe set, while the information preservation term evaluates the preservation of information structure with respect to pose changes. The distinct goals lead to different interpretations of relationships between descriptive vectors within the keyframe set.
Figure 3: Descriptors across poses. An example showing the descriptors, the Jacobian, and the transformed descriptors for two window keyframe sets. The Jacobian indicates how the features of the descriptors change between keyframes, while the transformed descriptors demonstrate how these same features appear after transformation to the principal components of that set.
III-C Sliding Window Optimization
To manage the computational complexity of optimizing a large keyframe set and proactively selecting keyframes for future queries, we introduce a sliding window optimization method. This approach continuously optimizes keyframe sets $\mathbf{K}_{t}$ over the mission duration $T$, approximating the optimal keyframe set as:
$\mathbf{K}_{\text{S}}^{*}\cong\bigcup_{t\,\in\,\left[0,T\right]}\mathbf{K}_{t}^{*},$ (12)
where $\mathbf{K}_{t}^{*}$ is the minimal subset of $\mathbf{K}_{t}$ that retains maximum information. To compute this subset, we use the redundancy and information preservation terms defined earlier.
The process begins by initializing a window keyframe set $\mathbf{K}_{t}$ containing $N$ keyframes. The time step $t$ advances once $N$ new keyframes are available, and the optimization converges to the optimal window keyframe set $\mathbf{K}_{t}^{*}$. To find this optimum, we generate all possible subsets of $\mathbf{K}_{t}$, forming the power set $\mathbb{P}(\mathbf{K}_{t})$ with cardinality $|\mathbb{P}(\mathbf{K}_{t})|=2^{N}$. To reduce computational complexity, we apply constraints on the power set, retaining only subsets that satisfy minimum and maximum distances between consecutive poses, as per Eq. (8), reducing the power set by 5–10 times. The constrained power set can be denoted as:
¯$\bar{\mathbb{P}}(\mathbf{K}_{t})=\big{\{}\mathbf{K}_{t}^{\mathbb{S}}\in\mathbb{P}(\mathbf{K}_{t})|\delta_{l}\leqslant\|\mathbf{x}_{i}-\mathbf{x}_{i+1}\|_{2}\leqslant\delta_{u},\forall\mathbf{k}\in\mathbf{K}_{t}^{\mathbb{S}}\big{\}},$ (13)
where $\delta_{l}$ and $\delta_{u}$ are the lower and upper limits for the distance between poses. The window optimization problem can be formulated as:
$\mathbf{K}_{t}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{K}_{t}^{\mathbb{S}}}\>\>\big{(}{\rho_{\tau}\,(\mathbf{K}_{t}^{\mathbb{S}})+\alpha}\big{)}\big{/}\big{(}{\pi_{\tau}(\mathbf{K}_{t}^{\mathbb{S}})}-\beta\big{)}$ (14)
$\operatorname*{where}\>\>\alpha,\,\beta>0\>\operatorname*{and}\>\mathbf{K}_{t}^{\mathbb{S}}\in\bar{\mathbb{P}}(\mathbf{K}_{t}).$ (15)
The minimization problem is solved through an exhaustive search, evaluating each subset’s information matrix $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ and quantifying its effectiveness using the information preservation and redundancy terms, selecting the subset with the best combined score.
IV Experimental Setup and Results
TABLE I: Translational (t) and Rotational (rot) Absolute Trajectory Error Improvement after PGO and the False Positive Rate (FPR)
Datasets KITTI Odometry (Seq. 00 / 06 ) MulRan (DCC / KAIST) Apollo-SouthBay (SanJose / Columbia)
Metrics t. ATE [%] rot. ATE [%] FPR [%] t. ATE [%] rot. ATE [%] FPR [%] t. ATE [%] rot. ATE [%] FPR [%]
All Samples 06.9 /-78.6 -06.1 /-51.1 01.4 /-05.6 29.8 / 67.4 24.9 / 59.5 08.0 / 21.2 96.2 / 98.1 95.3 / 97.5 30.7 / 07.5
Const. 1m 12.2 /-78.7 -12.3 /-55.3 02.8 /-07.0 59.3 / 87.1 56.4 / 71.0 20.2 / 23.8 85.8 / 97.1 84.1 / 97.2 25.1 / 04.1
Const. 2m 10.5 /-44.9 -06.5 /-23.7 01.3 /-03.1 70.0 / 80.9 68.4 / 73.7 07.6 / 11.6 06.8 / 75.9 06.8 / 77.2 25.3 / 02.3
Const. 3m 05.1 /-22.6 -03.9 /-13.8 03.0 /-09.5 63.6 / 87.7 64.6 / 79.1 04.4 / 07.7 00.0 / 83.6 00.0 / 84.4 16.6 / 02.4
Entropy 09.6 /-24.1 -03.8 /-09.1 01.5 /-05.4 69.2 / 68.7 69.5 / 63.1 16.4 / 14.6 77.6 / 69.4 76.2 / 72.6 21.0 / 03.1
Spacious. 04.4 /-35.5 -02.0 /-39.1 02.3 /-06.9 68.4 / 68.6 68.1 / 63.1 08.5 / 16.1 03.5 / 67.9 03.5 / 71.7 24.1 / 04.9
MSA (Ours) 11.1 /-78.8 -09.7 /-55.5 01.1 /-04.9 70.1 / 91.4 69.7 / 77.8 08.2 / 11.4 88.7 / 84.0 88.3 / 85.1 21.5 / 01.6
We evaluate our approach using three datasets: KITTI Odometry [34], MulRan [35], and Apollo-SouthBay [36], which cover diverse environments such as urban areas, rural landscapes, and complex structures like bridges and tunnels. Each dataset provides 3D LiDAR scans and GNSS-based ground truth poses, with MulRan being particularly challenging due to obstructed LiDAR views. For descriptor extraction, we use OverlapTransformer (OT) [24] with a $1\times 256$ feature vector. In [30], we demonstrate the generalizability of our method to other descriptors, like Scan Context [25]. Comparisons are made with constant intervals of 1, 2, and 3 meters [14, 17, 18, 15, 16], as well as with adaptive sampling based on LiDAR spaciousness [19] and entropy [21]. All experiments were conducted on a 14th Gen Intel Core i9-14900K with 128GB DDR5 RAM. The parameters of MSA are set to the defaults of $\alpha=\beta=1$ and $N=10$. The window optimization is solved, on average, in 14.4ms, making it suitable for real-time sampling.
IV-A Quantitative and Qualitative Analysis
For the batch experiments in Table I, the following pipeline is used: keyframes are sampled for all methods, and a pose graph is constructed using GTSAM [6] with odometry provided by KISS-ICP [37]. Potential loop closure matches are identified using OT descriptors, with a similarity threshold of 0.8. Ground truth data are used to classify candidates as true or false positives, based on a 1-meter radius. True positives are passed to small_gicp [38] for relative transformation estimation, with matches being verified if the registration residual is below 0.3 meters. Verified loop closures are then added as edges to the pose graph, which is optimized using the Levenberg-Marquardt algorithm.
Figure 4: Box plot comparisons. Translational (t) and Rotational (R) Absolute Trajectory Error (ATE) difference of every method compared to the baseline of All Samples, as well as the Memory Allocation in Gigabytes (GB) and the total Execution Time in seconds (s).
Figure 5: Trajectory comparisons. Comparison of the ground truth and KISS-ICP initial poses, and the sampled poses after the pose graph optimization for the proposed Minimal Subset Approach (MSA) and the entropy-based approach on the KAIST sequence of the MulRan dataset and the SanJoseDowntown sequence of the Apollo-SouthBay dataset.
Table I summarizes the results for all datasets and methods, where bold values represent the best performing adaptive method and underlined values indicate the best performing constant interval. The metrics include the Absolute Trajectory Error (ATE) for both translation and rotation, presented as the percent improvement in the KISS-ICP trajectory after pose graph optimization with each method. Additionally, the False Positive Ratio (FPR) for descriptor matching is provided, as sampling affects place recognition performance [30]. The results show that the proposed MSA consistently achieves the best performance, while constant intervals show varying results across different algorithms. In Fig. 4, the data from Table I are compiled into box plots, showing the deviation of each method’s ATE performance compared to using all samples, along with variations in memory usage and total execution time of the whole pipeline. The proposed approach maintains overall improvements with minimal performance loss in both translation and rotation, while significantly reducing memory usage and providing reasonable execution times. Although other methods reduce memory overhead and processing time, their performance varies considerably, lacking robustness. Finally, Figs. 1 and 5 visually compares ground truth poses, raw KISS-ICP poses, and the corrected trajectories for MSA and the entropy-based method after optimization. For all illustrated trajectories, the proposed approach consistently achieves lower translation and rotation RPE than the other methods, while the segments containing loop closures remain very close to the ground truth, demonstrating the approach’s ability to retain crucial keyframes.
IV-B Large-Scale Evaluation
TABLE II: Results for the Large-Scale Evaluation
Metrics t. ATE rot. ATE FPR MEM. TIME
All Samples nan nan nan > 128GB nan
Const. 1m 98.5% 98.4% 12% 90.9GB 7340s
Const. 2m 98.1% 97.3% 17% 47.2GB 2264s
Const. 3m 97.9% 97.0% 28% 35.4GB 1515s
Entropy 95.5% 93.4% 15% 51.0GB 4945s
Spacious. 94.9% 97.5% 24% 38.3GB 1752s
MSA (Ours) 98.3% 94.1% 12% 41.2GB 1951s
Figure 6: Trajectory comparisons. Comparison of the ground truth and KISS-ICP initial poses, and the sampled poses after the pose graph optimization for the proposed Minimal Subset Approach (MSA) and the entropy-based approach on the SunnyvaleBigloop sequence of the Apollo-SouthBay dataset where the color gradient denotes the RPE. A grid block corresponds to 0.5x0.5 km2
For the second part, we conduct a more realistic online experiment using the SunnyvaleBigloop sequence from Apollo-SouthBay, covering over 100 kilometers. Poses are processed sequentially, simulating real-time operation. Keyframes are sampled online using the previously mentioned methods, appended to the pose graph, and optimized at each step with iSAM2 [8]. Simultaneously, descriptors are queried against past keyframes to detect potential loop closures, which are verified with small_gicp and integrated into the graph. This online approach provides insights into how well each sampling method handles large-scale missions. Table II presents the translation and rotation ATE, FPR, total allocated memory, and processing time. Notably, when using all samples, the system runs out of memory approximately 70% into the sequence, so no metrics are provided. The remaining sampling methods reduce memory usage while offering excellent performance. This is likely because the mission revisits the same locations multiple times, allowing frequent loop closure edges that prevent significant accumulated drift. In Fig. 6, the optimized trajectory of the entropy-based method is compared to MSA, with the color gradient indicating the translational RPE, showcasing the better performance while maintaining lower memory allocation and false positive ratio. As a final step, Figs. 7 and 8 provide further insights into execution time and memory allocation. Fig. 7 shows memory usage over time, illustrating how the sampling methods scale more efficiently compared to retaining all samples or using a constant 1-meter interval. Fig. 8 presents the moving average of loop closure detection time (left) and incremental pose graph optimization time (right). The top figures show constant interval methods, while the bottom shows adaptive methods. The proposed approach scales similarly to other adaptive methods while offering more robust performance, as demonstrated earlier. The shaded areas represent the minimum and maximum deviations, with the moving average trend indicating the challenge of increasing processing time in large-scale missions. While query time exceeds 1 second per new sample, and the incremental pose graph optimization reaches up to 0.15s for the constant 1-meter sampling method, MSA retains query times of less than 0.5s and incremental optimization times of less than 25ms, maintaining the real-time performance of the system throughout the large-scale mission.
Figure 7: Memory allocation. Memory against mission time for the different sampling methods. Retaining all samples exceeds system memory of 128GB, demonstrating the need for efficient and scalable algorithms.
Figure 8: Processing times. The top row shows constant interval methods, while the bottom row shows adaptive methods. The left column represents loop closure detection time, and the right column represents incremental pose graph optimization convergence time. All times in seconds (s).
V Conclusions
In conclusion, as the robotics community advances towards larger-scale and long-term missions, addressing system limitations with efficient and scalable solutions becomes crucial. This article introduced the Minimal Subset Approach (MSA) to address the combined challenges of loop closure detection and pose graph optimization by reducing redundant samples while maintaining robust performance. Experimental evaluations across multiple datasets show that MSA outperforms in terms of FPR, as well as achieving superior ATE and RPE after pose graph optimization. Unlike other adaptive sampling methods, MSA requires minimal to no parameter tuning, while using a fraction of the system’s memory and maintaining comparable computational times.